Minutes of the Tracker Meeting
May 2, 2002
Tom reported that Alcomp proposes to deliver the flight CC material from August 02 through January 03 in several lots. We assume that the vendor who will machine the closeouts, however, will want to do all of them in one lot. The schedule presently indicates that the closeouts should be completed in April, giving about 3 months to machine them and install inserts.
Tom asked Erik whether all the testing proposed by Alcomp is necessary? Erik said that the density measurement indicates thermal properties but not necessarily mechanical. In any case, the testing should involve just a limited number of witness samples.
Sandro said that he received the box for storing trays from Bari and needs a drawing of the connector savers. Robert will carry an actual connector saver to Italy next week.
The last of the EM closeout material should be at SLAC Friday morning, in 2 shipments. Ossie will ship it immediately to Pisa.
Erik said that GMSI should have completed the EM closeout piece machining yesterday. They are now starting on bottom tray modifications for the new vibration experiment at Ames (this is the terminology that should be used henceforth in reference to this test, by directive from Project Management). Ossie will confirm today with GMSI. When they arrive at SLAC, he will ship them directly to Pisa.
Drawings for the insert tool were sent to Plyform. We will review them there on Tuesday.
The tray panel assembly tools are back at Plyform to correct the dimensional problem. Drawings by BJ are ready for the top-bottom tray tooling to be checked in Pisa.
Luisella is starting on the Tungsten assembly tool drawings.
Screws arrived yesterday (enough for the entire flight tracker). Torque tests look good. Ossie needs to do some acceptance testing: thread gauging and tests to destruction. They took 120 inch oz with no problem, while the specification is 110 inch oz.
Sandro requested that UCSC start putting together a mini-MCM system for Pisa. We decided to do this with f-version chips. It will be useful for learning to use the readout system but not for connecting to a detector ladder. UCSC is working on setting up a mini-MCM to use with a detector ladder using the d-version chips. Those chips have some bugs with loading the configuration register that can be worked around, and they have a problem with feedback caused by the internal preamp test pads, which can be corrected by laser cutting (a bit of a chore for 384 channels). This is the only means of verifying the detector system performance while waiting for the next submission.
Robert reported that 1 version of the front-end chip (the DC coupled version) is ready for submission. The AC coupled version layout is still being worked on (delayed by work on the calorimeter chips). The GTRC layout was done, but UCSC found another bug. That was understood and fixed in the design yesterday. Submission still looks a few days away.
Steve reported that the 2nd half of the test program on prototype trays was completed and written up. It will be on the web site in 1 week, after some internal review at Hytec. The 1st half of the tests is already on the web site. The basic conclusion is that there is no damage to trays or payload. The tests include dynamic, thermal, and vacuum.
Erik reported that the vibration data has been analyzed. There is now good correlation with the FEM. Lots of time went into understanding tray modes in the vertical axis test. The data frequencies are higher than in the model. Franz has concluded that there is a stiffer connection from wall to tray in the actual tower than in the model. Also, individual trays vibrating alone dont show up in the data. Instead, one sees all trays vibrating together in phase in two groups: heavy and light. There is a 180 phase between the heavy and standard trays but within each group they vibrate in phase. Erik is writing it up. He said that NASA sent the magnitude only when sending the data, not magnitude and phase. Because of that, he cannot get the displacements from the data. Ossie said that they will take ZIP disks next time to get the data on the spot.
Erik said that he has the bottom tray attachment test running from +50C to 30C in. After 10 cycles there is no evidence of damage. The strain data look consistent. He plans to continue to 120 cycles or until out of N2 or until the stuff needs to go back to SLAC. Ossie said that he is a week away from needing the stuff.